# Exploiting Sparsity for Long Context Inference: Million Token Contexts on Commodity GPUs

### v1

@@ 4.4 Layer and Task Dependence on k
We also examine the dependence of k on different layers. Recent works have shown that later layers of a transformer network tend to be less crucial to the computation than earlier ones (Gromov et al., 2024). These results motivated us to study whether the same value of k was necessary across all layers; the attention distributions of certain layers be well approximated with less k than others.
Figure 8 shows our study of k variance across layers. We vary k and plot the RULER scores as before, but we allocate k across layers in two different ways. In the uniform strategy each layer gets the same amount of k. In the adaptive strategy, we linearly increase k from the first to the last layer. At each point, the total k budget is the same: the sum across all layers of the k values used is equal. We are able to gain a non-trivial performance boost by changing our strategy of k.
@@ 5 Related Work
The problem of making transformer model inference more efficient has been approached from many angles, and we make a brief overview of some of the relevant works below.
@@ 6 Conclusion
We have demonstrated the capability of a top-k attention mechanism to operate at the million token scale on a single GPU. In addition, our investigation of attention distributions across layers points to future directions where the choice of $k$ can be adapted across layers. We achieve sublinear complexity and evaluate at over 95% accuracy on common benchmarks while using only 1% of the context length on average in the attention block. This exploitation of attention sparsity opens up new directions for efficient and viable solutions to long context inference in language models.
@@ A.1 Distribution of Attention Scores
In addition to looking at the attention distributions across tasks, we investigate if there are any systematic trends in the attention sparsity across layers of a model. Figure 9) shows the attention entropy for RULER subtasks when plotted by layer, and figure 3 shows the number of scores required to cover 75% of the full attention distribution, with a histogram plotting this value over hundreds of samples of Wikipedia text. Note that in both figures, the first layer clearly stands out as having the least concentrated attention distributions.

### v2

@@ 4.4 Exploring Non-Uniform Choices of k
Layer-Wise Settings for k
Recent work has shown that later layers of a transformer network tend to be less crucial to the computation than earlier ones (Gromov et al., 2024). Our method naturally admits a dimension of flexibility as to where efficiency gains are extracted. As some layers may need a larger k budget than others to accurately capture their attention distributions, we explore allowing the value of k to vary across layers.
In Figure 9 we vary k and plot the RULER scores as before, but we allocate our budget for k across layers in two different ways. In the uniform strategy, each layer gets the same value of k. In the adaptive strategy, we linearly increase k from the first to the last layer. In each experiment (point on chart), the total k budget remains fixed, e.g. the sum of k over all layers is always the same. We find we are able to gain non-trivial performance boosts simply by changing our strategy for how set k for each layer given a fixed computational budget.
@@ 5 Related Work
The problem of making transformer model inference more efficient has been studied from many angles. We briefly survey some of the relevant work below.
@@ 6 Conclusion
In this work we demonstrate the capability of a top-k attention mechanism to operate at the million token scale on a single GPU. We achieve sublinear complexity and evaluate at over 95% of dense attention accuracy on common benchmarks while using only 2% of the context length on average in the attention block. This exploitation of attention sparsity opens up new directions for efficient and viable solutions to long context inference in language models. Our investigation of attention distributions across layers points to future variations of our method that smoothly adapt the top-k method across tasks and layers of a given model suggesting our top-k approach can be used to achieve optimal compute-performance tradeoffs given a specific deployment context.
@@ A.1 Distribution of Attention Scores
In addition to looking at the attention distributions across tasks, we investigate if there are any systematic trends in the attention sparsity across layers of a model. Figure 10 shows the attention entropy for RULER subtasks when plotted by layer, and Figure 3 shows the number of scores required to cover 75% of the full attention distribution, with a histogram plotting this value over hundreds of samples of Wikipedia text. Note that in both figures, the first layer clearly stands out as having the least concentrated attention distributions.